OpenStudy (drewbooks):

Find the GCF: 5x^5y^2 + 15x^2y^3 + x^3y^5

OpenStudy (drewbooks):

I believe B is correct, but I'm not certain. @Er.Mohd.AMIR A) xy B) x^2y^2 C) 5xy D 5x^2y^2

OpenStudy (er.mohd.amir):

again factor all terms and take common . 5x^5y^2=5*x*x*x*x*x*y*y

OpenStudy (drewbooks):

15x2y3 = 15*x*x*y*y*y x^3y^5 = x*x*x*y*y*y*y*y Right? So, would that not make x^2y^2 correct?

OpenStudy (er.mohd.amir):

x^2y^2 is right because no number in third one only x^2 and y^2 is in all terms

OpenStudy (drewbooks):

Got it, thank you.